OpenStudy (anonymous):

Need some algebra 2 help, please! Would be highly appreciated

OpenStudy (anonymous):

1.Create a rational expression to be your game piece. You may choose from the list of factors below or make your own. There must be a variable term in both the numerator and denominator. •(5x) •(2x) •(x + 4) •(x – 5) •(2x + 1) •(3x + 1) I"m not sure as to what rational I should make. I was thinking of using the (2x)

OpenStudy (anonymous):

@agent0smith @jim_thompson5910 Can one of you help please

jimthompson5910 (jim_thompson5910):

Simply pick two expressions and divide them. That's how you'll create a rational expression.

OpenStudy (agent0smith):

Just choose a couple... I pick (2x)/(3x + 1)

OpenStudy (anonymous):

Well, the trick is that it's like a "board game" so I need to flip a coin, and whatevr side it lands on I have to solve it using the expression i guess?

jimthompson5910 (jim_thompson5910):

seems like you start with 0 then if you flip heads on the first coin flip, then you add on \(\Large \frac{x+1}{2x}\) to 0 or if you flip tails on the first coin flip, then you add on \(\Large \frac{x-1}{3x}\) to 0

OpenStudy (anonymous):

I think we have to create one to add it to. Like we start out with our created "game piece" and if i flip tails then I have to solve my rational experession with the other.

jimthompson5910 (jim_thompson5910):

oh i see, that works too

OpenStudy (anonymous):

So we could use Agent's expression? Like I flipped heads so it would be\[ \frac{ 2x }{ 3x+1 } + \frac{ x+1 }{ 2x }\]

jimthompson5910 (jim_thompson5910):

yes, then you would add and simplify

OpenStudy (anonymous):

Could you explain how to for me? So I can do it for the next 4 questions?

jimthompson5910 (jim_thompson5910):

in general, how do you add fractions let's say you wanted to add \(\Large \frac{1}{2}+\frac{1}{3}\) how would you do it?

OpenStudy (anonymous):

Find a common denominator?

jimthompson5910 (jim_thompson5910):

which is what?

OpenStudy (anonymous):

Six. Times the 2 by 3 and times the 3 by 2, right?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

in the first fraction, you multiplied the denominator by 3 so you have to do the same to the numerator to balance it out so \[\Large \frac{1}{2} = \frac{1*3}{2*3} = \frac{3}{6}\] see how I'm doing this?

OpenStudy (anonymous):

Yes, I'm following along :)

jimthompson5910 (jim_thompson5910):

Similarly, \[\Large \frac{1}{3} = \frac{1*2}{3*2} = \frac{2}{6}\]

jimthompson5910 (jim_thompson5910):

which means \[\Large \frac{1}{2}+\frac{1}{3}\] is the same as \[\Large \frac{3}{6}+\frac{2}{6}\]

jimthompson5910 (jim_thompson5910):

now that the denominators are the same, you can add the numerators and place them over the common denominator

jimthompson5910 (jim_thompson5910):

so you apply these basic steps to add rational expressions

OpenStudy (anonymous):

So since my denominators have a 3 and a 2, would it be 6 again for this as well? It's confusing when it haves the 3x+1

jimthompson5910 (jim_thompson5910):

in that last example, the denominators were 2 and 3 the LCD is 6 since 2*3 = 6 (and the GCF of 2 and 3 is 1)

jimthompson5910 (jim_thompson5910):

so if there is no common factor between the denominators (other than 1), you just multiply the denominators to get the LCD

OpenStudy (anonymous):

So I would muiltiply 3x+1 + 2y? sorry if i'm frustrating you i'm just really confused

jimthompson5910 (jim_thompson5910):

the denominators in this case are 3x+1 and 2x multiply them to get 2x(3x+1) you can leave it factored or you can distribute

jimthompson5910 (jim_thompson5910):

so the LCD is 2x(3x+1)

jimthompson5910 (jim_thompson5910):

the goal is to get every denominator equal to the LCD

OpenStudy (anonymous):

So it would now be \[\frac{ 2x }{2x(3x+1)} + \frac{ x+1 }{ 2x(3x+1) }\]

jimthompson5910 (jim_thompson5910):

you have the denominators correct, but the numerators are incorrect

jimthompson5910 (jim_thompson5910):

in the first fraction, you need to multiply top and bottom by 2x to get \[\Large \frac{ 2x }{ 3x+1 }\] \[\Large \frac{ 2x*2x }{ 2x(3x+1) }\] \[\Large \frac{ 2*2*x*x }{ 2x(3x+1) }\] \[\Large \frac{ 4x^2 }{ 2x(3x+1) }\] You have to multiply top and bottom of the fraction to balance it out

OpenStudy (anonymous):

So since I rolled heads and we found the LCD would this be the answer for this question. Explain to the game master how to add your rational expression to the one on the correct space. Or do we still have to add together now?

jimthompson5910 (jim_thompson5910):

no you need to add the rational expressions

OpenStudy (anonymous):

Okay, so how could we do that?

jimthompson5910 (jim_thompson5910):

you get each fraction to get the LCD using the methods shown above

jimthompson5910 (jim_thompson5910):

once the fractions have the same denominator, you can add the numerators and place them over the LCD

OpenStudy (anonymous):

So 4x^2 + 4x2 would be 8x^2 or 32x?

jimthompson5910 (jim_thompson5910):

for the second fraction, you have to multiply top and bottom by 3x+1

OpenStudy (anonymous):

What do you mean

jimthompson5910 (jim_thompson5910):

\[\Large \frac{ x+1 }{ 2x }\] \[\Large \frac{ (3x+1)(x+1) }{ (3x+1)2x }\] \[\Large \frac{ 3x(x+1)+1(x+1) }{ 2x(3x+1) }\] I'll let you finish simplifying

OpenStudy (anonymous):

I'm stumped man, I'm trying to figure it out I promise it's just super confusing. So would I do \[\frac{ 3x(x+1 + 1 (x-1) }{ 2x (3x +1) }\] And then \[4x + 1 (x+1) \] For the numerator? Sorry if i'm totally failing

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

3x(x+1) + 1(x+1) 3x*x + 3x*1 + 1*x + 1*1 3x^2 + 3x + 1x + 1 3x^2 + 4x + 1

jimthompson5910 (jim_thompson5910):

so, \[\Large \frac{ 2x }{ 3x+1 } + \frac{ x+1 }{ 2x }\] is the same as \[\Large \frac{ 4x^2 }{ 2x(3x+1) } + \frac{ 3x^2+4x+1 }{ 2x(3x+1) }\]

OpenStudy (anonymous):

So now how do we solve?

jimthompson5910 (jim_thompson5910):

the denominators are now equal, so you can add the numerators and place that result over the LCD

OpenStudy (anonymous):

Are we doing a combining like terms kinda thing?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

so \[\frac{ 3^{x2 } + 4x +1}{2x(3x+1 }\]

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

:/ sorry

jimthompson5910 (jim_thompson5910):

3x^2 + 4x^2 is ????

OpenStudy (anonymous):

7x^2

jimthompson5910 (jim_thompson5910):

\[\Large \frac{ 2x }{ 3x+1 } + \frac{ x+1 }{ 2x }\] \[\Large \frac{ 4x^2 }{ 2x(3x+1) } + \frac{ 3x^2+4x+1 }{ 2x(3x+1) }\] \[\Large \frac{ 4x^2 + 3x^2+4x+1 }{ 2x(3x+1) }\] \[\Large \frac{ 7x^2+4x+1 }{ 2x(3x+1) }\]

OpenStudy (anonymous):

Thank you! Sorry it took me a long time

jimthompson5910 (jim_thompson5910):

that's ok, and yw